The average marks in the mathematics of a sample of 100 students was 51 with a S.D. of 6 marks .Could this have been a random sample from the population with average marks 50.


1
Expert's answer
2021-05-07T14:23:02-0400

Solution:

Here "n=100, \bar{x}=51, s=6, \mu=50" ; "\sigma" is unknown.

"\mathrm{H}_{0}" : The sample is drawn from a population with mean "50, \mu=50"

"\mathrm{H}_{1}: \mu \neq 50"

Under "\mathrm{H}_{0}, z=\frac{\bar{x}-\mu}{s / \sqrt{n}}=\frac{51-50}{6 / \sqrt{100}}=\frac{10}{6}=1.6666" .

Conclusion. Since "|z|=1.666<1.96, z_{\alpha}" the significant value of z at 5% level of significance, "\mathrm{H}_{0}" is accepted, i.e., the sample is drawn from the population with mean 50 .